# Cost-Effective, Low Latency Vector Search with Azure Cosmos DB

### v1

@@ Abstract.
Vector indexing enables semantic search over diverse corpora and has become an important interface to databases for both users and AI agents. Efficient vector search requires deep optimizations in database systems. This has motivated a new class of specialized vector databases that optimize for vector search quality and cost. Instead, we argue that a scalable, high-performance, and cost-efficient vector search system can be built inside a cloud-native operational database like Azure Cosmos DB while leveraging the benefits of a distributed database such as high availability, durability, and scale. We do this by deeply integrating DiskANN, a state-of-the-art vector indexing library, inside Azure Cosmos DB NoSQL. This system uses a single vector index per partition stored in existing index trees, and kept in sync with underlying data. It supports ¡ 20ms query latency over an index spanning 10 million of vectors, has stable recall over updates, and offers nearly 15×15$\times$15 × and 41×41$\times$41 × lower query cost compared to Zilliz and Pinecone serverless enterprise products. It also scales out to billions of vectors via automatic partitioning. This convergent design presents a point in favor of integrating vector indices into operational databases in the context of recent debates on specialized vector databases, and offers a template for vector indexing in other databases.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
@@ 1. Introduction
Advances in deep learning have made it possible to embed data as vectors in high-dimensional vector spaces so that the distance between vectors captures various notions of semantic similarity. This opens up a new interface for users as well as AI models/agents to interact with large information stores based on semantic and contextual relevance, rather than literal matches or structured queries. Therefore, efficient search in vector spaces has become a critical requirement for information retrieval systems. Already, vector search is a central component in industrial scale retrieval (web and document search) and recommendation systems. In databases, especially document databases, augmenting existing workloads with vector representations is becoming commonplace.
A new class of scenarios requiring vector search on operational data modeling e-commerce, document retrieval, conversational histories, and AI agent interaction patterns, is rapidly growing. This has motivated a new class of specialized vector databases that optimize primarily for vector search performance. However, this pattern forces the replication of data between a primary operational database and a secondary vector database which can cause data divergence, and increased cost and complexity for the user. This also may not provide the operational resilience that developers expect.
@@
  • Elastic scaling to billions of thousand+ dimensional vectors.
@@
DiskANN is a suite of vector indexing algorithms (Krishnaswamy et al., 2025; Subramanya et al., 2019; Singh et al., 2021) designed for such constraints. It derives quantized representations of vectors which can be much smaller than the original vectors, and supports updates and queries to the index mostly via quantized vectors. To perform an index update or query, full-precision vectors stored in the index are accessed 50×50$\times$50 × less frequently compared to quantized vectors. This allows the system to provide high performance even when most of the index is stored on SSDs. DiskANN has been widely deployed in several extreme scale Microsoft semantic indices used for web search, advertisements, Microsoft 365 search, Co-pilots as well as on edge devices (Pavan Davuluri, [n.d.]).
The DiskANN library (Krishnaswamy et al., 2025) has previously been designed to control the layout of index terms either in memory or in SSD, akin to other monolithic systems  (Johnson et al., 2017; HNS, 2019). While databases such as SingleStore (Chen et al., 2024) and Elastic (Pentreath et al., 2017; Tibshirani, 2019) use vector indexing libraries in a loosely coupled way to produce a separate index for each immutable data segment, we do not use such a design due to several drawbacks: (a) each query has to fan out to numerous segments which limits query efficiency, (b) regular rebuilds of vector indices are needed as segments are merged or consolidated, which consumes significant compute, memory and causes serious latency spikes for queries (Trent, 2023, 2024), (c) either large cold start latencies occur while loading large vector indices into memory, or high expense is incurred from having to hold the index in memory.
Instead, for simplicity and robustness of the system, we store the terms representing the vector index on the Bw-Tree index in Cosmos DB. Bw-Tree supports high concurrency through latch free algorithms, provides fast random reads and writes in a tiered memory/SSD setting. To operate in this setting, we rewrote the DiskANN library to decouple the algorithmic logic from physical index layout. The new library provides supports updates and queries by manipulating or reading index terms (quantized vector and graph adjacency lists) stored external to the library. This leads to several structural advantages:
  • We can maintain and update just one vector index per replica, which can be as large as 50GB, enabling higher query efficiency via reduced fan-out. In fact, DiskANN’s query complexity scales logarithmically in the size of the index.
@@
  • The Bw-Tree caches index terms for hot partitions on demand, while collections that have not been used recently do not waste memory and can only be billed for storage.
  • The Bw-Tree has a long history of support in the current system, so its reuse for DiskANN indexing terms benefits from its established stability and ease of operation.
@@
This deep integration eliminates the need for a vector index outside an operational database, and instead composes sophisticated features from Cosmos DB and DiskANN. From the new DiskANN rewrite we inherit existing features – querying with limited memory – as well as new features developed for this integration such as index updates with limited memory, filter-aware search and paginated search. The latter increases the efficiency of hybrid queries with predicates and vector search. From Cosmos DB, we inherit flexible cost structures, resource governance, elasticity, and resiliency.
@@
  • Query cost that is nearly 15×15$\times$15 × and 41×41$\times$41 × lower than Zilliz and Pinecone enterprise-tier serverless vector databases respectively (for 10 million 768D vectors), while offering higher availability.
@@
  • Ingest offers stable recall over long update sequences, cost and performance comparable to other vector databases.
@@ 2.1. The DiskANN vector indexing library
The index consists of a graph over the vectors in the database, with one vertex representing each vector and directed edges connecting vertices. The search for a query $q$ uses a “greedy” approach, starting at a designated start point $s$, computing the distances from $q$ to each point in the out-edges $N_{\rm out}({s})$, and moving on to the nearest neighbor of $q$ among $N_{\rm out}({s})$. It continues this process of greedily visiting the closest neighbor to $q$ until it can no longer improve on the closest neighbor to $q$, at which point the search terminates. This algorithm is formally described in Algorithm 1, and can be naturally extended to return the top-$k$ neighbors by keeping a priority queue of the $k$ closest neighbors instead of hopping to the single closest neighbor each time. The accuracy, or recall k@k, of a search, is defined as how many of the $k$ results returned by a search are the true top-$k$ nearest neighbors. When clear from context, it is also used to refer to the average recall over a batch of searches.
DiskANN’s query complexity grows logarithmically with the size of the index (empirically observed (Subramanya et al., 2019; Manohar et al., 2024b)). It can work effectively with an almost entirely SSD-based index and limited memory, while providing performance parity with in-memory indices like ScaNN that consume an order of magnitude more memory. DiskANN is IO efficient – it can achieve 90% recall@10 on a billion-size SIFT dataset with as few as 50 random 4 KB reads to SSD. The same DiskANN index can also be loaded entirely into DRAM for scenarios requiring extreme throughput, where it outperforms graph-based methods such as HNSW (HNSW, [n.d.]) and partition-based methods such as IVF and ScaNN (Manohar et al., 2024a).
@@
      return [[[[closest $k$ points from $\mathcal{V}$; $\mathcal{V}$]]]]
@@
The DiskANN graph is built via repeated calls to the insertion algorithm, which is formally described in Algorithm 2. At a high level, the insert procedure generates candidates for insertion using a call to Algorithm 1, prunes the candidates down to respect the degree bound $R$, and then adds edges pointing to the newly inserted node to make it reachable. One of the main innovations behind DiskANN’s performance is the RobustPrune routine, shown in Algorithm 3, which is used to prune a vertex’s out-neighbors down to the degree bound $R$. At a high level, it removes an edge $(u,v)$ when $v$ is likely to be reachable via one of $u$’s other neighbors. Furthermore, it utilizes a scaling parameter $\alpha$ to prune more or less aggressively; in practice, the ability to scale $\alpha$ to prune less aggressively is consequential for performance (Subramanya et al., 2019; Singh et al., 2021). In some cases, it may also be necessary to replace the vector corresponding to a document identifier. We handle this by overwriting the original vector and invoking Algorithm 2 to re-insert the new vector. Any pre-existing edges pointing to the replaced point are cleaned up lazily via later calls to pruning.
@@
When a document is deleted, the corresponding vector and quantized vector are immediately removed from the index. Diskann removes updates the graph index terms to reflect the deletion using Algorithm 6, which is an adapatation of (Xu et al., 2025a). At a high level, the algorithm eagerly replaces the connections between non-deleted points that were maintained via a deleted point, ensuring stability of the index quality. A lightweight background process continuously removes any remaining edges pointing to the deleted point. Experiments show that the combination of in-place deletion and lightweight background consolidatipn is effective at keeping recall stable over long cycles of insertion and deletion.
@@
An additional algorithmic building block is the compression of the associated vector data. This allows data to be stored more compactly in expensive storage tiers (e.g., main memory), transmitted more efficiently across memory bus, and distance comparisons to be computed with fewer CPU cycles with little loss in accuracy. Both scalar quantization and product quantization are widely used on top of a vector index to increase search speeds. Scalar quantization maps each coordinate of the embedding to a smaller representation. For example, 32-bit floating point representations are easily rounded to the nearest 16-bit floating point with little loss of precision. Rounding to the nearest 8-bit integer or even to 1- or 2-bit representations is more at a coordinate level, but still preserves enough information overall to help the query navigate the index.
Product quantization (PQ) (Jegou et al., 2010) maps collections of coordinates to a few bytes by clustering data and mapping the data to the identity of the coordinate center. For many datasets, product quantization achieves a better compression than scalar quantization normalized for noise introduced in distance calculations. For example, PQ can compress OpenAI’s ada-003 embeddings (12KB) by 96x while retaining enough information to navigate the index. While PQ was specifically formulated for preserving Euclidean distances between embeddings, in practice it is also reasonable at preserving inner product distances.
@@ 2.2. The Cosmos DB system
Cosmos DB uses the simplicity of JSON and its lack of a schema specification. No assumptions are made about the documents stored in Cosmos DB and they can vary in schema. Cosmos DB operates directly at the level of JSON grammar, blurring the boundary between the structure and instance values of documents. This, in turn, allows the database to be ”schema-free”, enabling it to automatically index documents without requiring schema or secondary indices. For additional control, a custom indexing policy(Corporation, 2024b) can also be used to index specific properties using the ’path’ notation, that allows precise navigation to specific substructures in a JSON document. For example, the property path ’/employee/name’ represents the ’name’ node in the ’employee’ object. Cosmos DB, in addition to automatic indexing supported for the JSON type system, also has support for specialized indexes. This includes Spatial Indexing and more recently, indexing for vector and full-text search. For fast query and avoiding storage bloat associated with JSON text, Cosmos DB employs a performant custom binary encoding (cos, 2024) for JSON data.
@@
Clients define a logical partition key on a Collection (up to three levels of hierarchy are allowed in partition keys (Corporation, 2024a)). A Collection can thus span multiple physical partitions, with data hashed horizontally across partitions based on the logical partition key value. As clients adjust throughput and/or storage needs on their Collections, the elasticity component can scale out or scale back the compute and storage required through partition splits and merges.
@@
Cosmos DB service is deployed worldwide on clusters of machines each with dedicated local SSDs. The unit of deployment, called a federation (Figure 1), is an overlay network of machines, which can span one or more clusters. Each physical machine hosts replicas corresponding to various partitions for scaled out collections. Replicas corresponding to a single partitions are placed and load balanced across machines spanning different fault domains, upgrade domains and availability zones in the federation. Each replica hosts an instance of Cosmos DB database engine, which manages the JSON documents as well as the associated indices.
@@ 3. System Design
An improvement would be to map all vectors in the collection to a contiguous range in the Bw-Tree by prefixing the vectors with the collection id and the name of the vector index path (we call this the Flat index). This can be done easily in the user request path as part of the ”Document Analysis” during ingestion on the primary replica, and then replicated to all secondaries. A range scan in the Bw-Tree for the appropriate prefix would need many fewer random accesses for smaller vectors. This is still not a great option since vectors tend to be large, and storing them twice limits scale up.
A second improvement would be to compress the vectors via quantization (say using PQ) and store them in a contiguous range in the Bw-Tree. This significantly reduces the number of nodes to scan in the Bw-Tree, by up to 96×96$\times$96 × for OpenAI Ada v3 embeddings, for example. To answer a request for the top-$k$ nearest neighbors to query $q$, we would first find, say, $5k$ entries to the query in the quantized space. We could then look up the full precision vectors corresponding to each of the $5k$ candidates from Document Store, and compute the full fidelity distance to the query to identify the top-$k$ candidates. We refer to this as the Q-Flat (Quantized Flat) index. With an appropriate multiplier over how many extra candidates are retrieved in quantized space, this method can yield very high recall. For moderate sized collections, or for small tenants in a multi-tenant collection, this can be efficient. For instance, in a collection of $5,000$ vectors quantized to $96$ bytes, we need to touch only about $60$ 8KB-sized Bw-Tree logical leaf nodes to answer the query. The CPU time required for computing distances over the retrieved quantized vectors is under a millisecond.
However, an exhaustive scan in quantized space does not scale to larger replicas. Cosmos DB can fit over 10 million 768-dimensional floating point vectors in one replica, and we need a index that can answer queries by accessing fewer nodes in Bw-Tree. We use the $\rm DiskANN$ graph-structured index for this. With the $\rm DiskANN$ vector indexing policy in a collection, two additional indexing terms in the Bw-Tree are created: quantized vectors (akin to Q-Flat) and graph neighbor terms which represent out-neighbors of each vector in the index (see Fig. 2). In the rest of this section, we describe how we layer these index terms in Cosmos DB, and how $\rm DiskANN$ manipulates the terms to update and query the index.
@@ 3.1. Re-designing DiskANN for databases
$\rm DiskANN$ was previously written as a library that manages its own buffers and index layout, thus limiting its use inside a database. We rewrote it in Rust using the following principles so that it can used with a variety of systems including databases and key-value stores.
@@
The index layout is not controlled or even visible to the algorithms. The core of the library consists of methods that update and query the index by reading and updating index terms – quantized vectors, full precision vectors and neighbor list for vertices – via implementations of standardized Provider traits. The NeighborProvider trait, for example, defines the way the index would retrieve, append and overwrite the out-neighbors of a vector. These traits are implemented by the database which understands the best lay out and encoding/decoding for each kind of indexing terms. The database also owns the persistence and recovery of these terms.
@@
The term that the index needs may be immediately available in a memory buffer or may need to be retrieved from a slower storage device. Therefore the Provider traits allow for get methods to return either the actual data or a future that eventually returns the data so that the calling thread can be scheduled with other work meanwhile. This is encoded via the MaybeDone enumerated type in the snippet below:
@@
As a consequence, all update and query methods in the library are also asynchronous and need a runtime to manage threads and drive the futures to completion. We use the tokio runtime (tok, 2016).
@@
Our rewrite achieved these goals without compromising on performance compared to monolithic “in-memory” libraries such as (HNS, 2019; Simhadri et al., 2023). We can implement the Provider traits using a type backed by memory buffers for maximum performance – in fact, the new library is at least as fast as the the previous monolithic $\rm DiskANN$ library for all use cases it supported.

### v2

@@ Abstract.
Vector indexing enables semantic search over diverse corpora and has become an important interface to databases for both users and AI agents. Efficient vector search requires deep optimizations in database systems. This has motivated a new class of specialized vector databases that optimize for vector search quality and cost. Instead, we argue that a scalable, high-performance, and cost-efficient vector search system can be built inside a cloud-native operational database like Azure Cosmos DB while leveraging the benefits of a distributed database such as high availability, durability, and scale. We do this by deeply integrating DiskANN, a state-of-the-art vector indexing library, inside Azure Cosmos DB NoSQL. This system uses a single vector index per partition stored in existing index trees, and kept in sync with underlying data. It supports ¡ 20ms query latency over an index spanning 10 million vectors, has stable recall over updates, and offers approximately 43×43$\times$43 × and 12×12$\times$12 × lower query cost compared to Pinecone and Zilliz serverless enterprise products. It also scales out to billions of vectors via automatic partitioning. This convergent design presents a point in favor of integrating vector indices into operational databases in the context of recent debates on specialized vector databases, and offers a template for vector indexing in other databases.
PVLDB Reference Format:
PVLDB, 18(12): 5166 - 5183, 2025.
doi:10.14778/3750601.3750635 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 12 ISSN 2150-8097.
doi:10.14778/3750601.3750635
@@ 1. Introduction
Advances in deep learning have made it possible to embed data as vectors in high-dimensional vector spaces so that the distance between vectors captures various notions of semantic similarity. This opens up a new interface for users as well as AI models and agents to interact with large information stores based on semantic and contextual relevance, rather than literal matches or structured queries. Therefore, efficient search in vector spaces has become a critical requirement for information retrieval systems. Already, vector search is a central component in industrial scale retrieval (web and document search) and recommendation systems. In databases, especially document databases, augmenting existing workloads with vector representations is becoming commonplace.
A new class of scenarios requiring vector search on operational data modeling e-commerce, document retrieval, conversational histories, and AI agent interaction patterns, are rapidly growing. This has motivated a new class of specialized vector databases that optimize primarily for vector search performance. However, this pattern forces the replication of data between a primary operational database and a secondary vector database which can cause data divergence, and increased cost and complexity for the user. This also may not provide the operational resilience that developers expect.
@@
  • Elastic scaling to billions of vectors with thousands of dimensions.
@@
DiskANN is a suite of vector indexing algorithms (Krishnaswamy et al., 2025; Subramanya et al., 2019; Singh et al., 2021) designed for such constraints. It derives quantized representations of vectors which can be much smaller than the original vectors, and supports updates and queries to the index mostly via quantized vectors. To perform a query, full-precision vectors stored in the index are accessed 50×50$\times$50 × less frequently compared to quantized vectors. This allows the system to provide high performance even when most of the index is stored on SSDs. DiskANN has been widely deployed in several extreme scale: Microsoft semantic indices used for web search, advertisements, Microsoft 365 search, Co-pilots as well as on edge devices (Pavan Davuluri, 2024).
The DiskANN library (Krishnaswamy et al., 2025) has previously been designed to control the layout of index terms either in memory or in SSD, akin to other monolithic systems  (Johnson et al., 2017; et al., 2019). While databases such as SingleStore (Chen et al., 2024) and Elastic (Pentreath et al., 2017; Tibshirani, 2019) use vector indexing libraries in a loosely coupled way to produce a separate index for each immutable data segment, we do not use such a design due to several drawbacks: (a) each query has to fan out to numerous segments which limits query efficiency, (b) regular rebuilds of vector indices are needed as segments are merged or consolidated, which consumes significant compute, memory and causes serious latency spikes for queries (Trent, 2023, 2024), (c) either large cold start latencies occur while loading large vector indices into memory, or high expense is incurred from having to hold the index in memory.
Instead, for simplicity and robustness of the system, we store the terms representing the vector index on the Bw-Tree index in Cosmos DB. Bw-Tree supports concurrency through latch free algorithms, provides fast random reads and writes in a tiered memory/SSD setting. To enable this, we rewrote the DiskANN library to decouple the algorithmic logic from physical index layout. The new library provides supports updates and queries by manipulating or reading index terms (quantized vector and graph adjacency lists) stored external to the library.
This leads to several structural advantages:
  • We can maintain and update just one vector index per replica, which can be as large as 50GB, enabling higher query efficiency via reduced fan-out. In fact, DiskANN’s query complexity scales logarithmically (empirically) in the size of the index.
@@
  • The Bw-Tree caches index terms for hot partitions on demand, while cold collections do not consume memory.
@@
This deep integration eliminates the need for a vector index outside an operational database, and instead composes proven features from Cosmos DB and DiskANN. From Cosmos DB, we inherit flexible cost structures, resource governance, elasticity, and resiliency. By storing its indexing terms in Bw-Tree, which has a long history of support in Cosmos DB, DiskANN benefits from its established stability and ease of operation. From the new rewrite of DiskANN we inherit existing features – querying with limited memory – as well as new features developed for this integration such as index updates with limited memory, and filter-aware and paginated search for efficiently processing hybrid vector queries with predicates.
@@
  • Query cost that is nearly 12×12$\times$12 × and 43×43$\times$43 × lower than Zilliz and Pinecone enterprise-tier serverless vector databases respectively (for 10 million 768D vectors), while offering higher availability.
@@
  • Ingest offers stable recall over long update sequences, with cost and performance comparable to other vector databases.
@@ 2.1. The DiskANN vector indexing library
[…]
DiskANN’s query complexity grows logarithmically with the size of the index (empirically observed (Subramanya et al., 2019; Manohar et al., 2024b)). It can work effectively with an almost entirely SSD-based index and limited memory, while providing performance parity with in-memory indices like ScaNN that consume an order of magnitude more memory. DiskANN is IO efficient – it can achieve 90% recall@10 on a billion-size SIFT dataset with just 50 random 4 KB reads to SSD. The same DiskANN index can also be loaded entirely into DRAM for scenarios requiring extreme throughput, where it outperforms graph-based methods such as HNSW (HNSW, 2023) and partition-based methods such as IVF and ScaNN (Manohar et al., 2024a).
@@
 return [[[closest $k$ points from $\mathcal{V}$; $\mathcal{V}$]]]
@@
[…]
@@
When a document is deleted, the corresponding vector and quantized vector are immediately removed from the index. Diskann removes the graph index terms to reflect the deletion using Algorithm 6 in the Appendix, which is an adaptation of (Xu et al., 2025). At a high level, the algorithm replaces the critical connections to the deleted point, ensuring stability of the index quality. A lightweight background process continuously removes any remaining edges pointing to the deleted point. Experiments show (Xu et al., 2025) that the combination of in-place deletion and lightweight background consolidation is effective at keeping recall stable over long cycles of insertion and deletion.
@@
An additional algorithmic building block is quantization to compress the vectors. This allows them to be stored more compactly in expensive storage tiers (e.g., main memory), transmitted more efficiently across memory bus, and distance comparisons to be computed with fewer CPU cycles with little loss in accuracy. One popular method is scalar quantization which maps each coordinate of the embedding to a smaller representation. For example, 32-bit floating point representations are easily rounded to the nearest 16-bit floating point with little loss of precision. Rounding to the nearest 1- or 2-bit representations is more lossy at a coordinate level, but could still preserve enough information overall for the query to navigate the index.
Product quantization (PQ) (Jegou et al., 2010) maps collections of coordinates to a few bytes by clustering data and mapping the data to the identity of the coordinate center. For many datasets, product quantization achieves better compression than scalar quantization normalized for errors introduced in distance calculations. For example, PQ can compress OpenAI’s text-3-large embeddings (12KB) by 96x while retaining enough information to navigate the index. Although PQ was formulated for preserving euclidean distances between vectors, in practice it also preserves inner product distances reasonably well.
@@ 2.2. The Cosmos DB system
Cosmos DB uses the simplicity of JSON and its lack of a schema specification. No assumptions are made about the documents stored and they can vary in schema. Cosmos DB operates directly at the level of JSON grammar, blurring the boundary between the structure and instance values of documents. This, in turn, allows the database to be ”schema-free”, enabling it to automatically index documents without requiring schema or secondary indices. For additional control, a custom indexing policy(Corporation, 2024b) can also be used to index specific properties using the ’path’ notation, that allows precise navigation to specific substructures in a JSON document. For example, the property path ’/employee/name’ represents the ’name’ node in the ’employee’ object. Cosmos DB, in addition to automatic indexing supported for the JSON type system, also supports specialized indexes such as spatial indexing, and more recently, indexing for vector and full-text search. For faster query and to avoid storage bloat associated with JSON text, Cosmos DB employs a custom binary encoding (Souza, 2024) for JSON data.
@@
Clients define a logical partition key on a Collection (up to three levels of hierarchy are allowed in partition keys (Corporation, 2024a)). A Collection can thus span multiple physical partitions, with data hashed horizontally across partitions based on the logical partition key value. As clients adjust throughput and/or storage needs, the compute and storage required for collections are scaled out or scale back through partition splits and merges.
@@
Cosmos DB service is deployed worldwide on clusters of machines each with dedicated local SSDs. The unit of deployment, called a federation (Figure 1), is an overlay network of machines, which can span one or more clusters. Each physical machine hosts replicas corresponding to various partitions for scaled out collections. Replicas corresponding to a single partition are placed and load balanced across machines spanning different fault domains, upgrade domains and availability zones in the federation. Each replica hosts an instance of Cosmos DB database engine, which manages the JSON documents as well as the associated indices.
@@ 3. System Design
An improvement would be to map all vectors in the collection to a contiguous range in the Bw-Tree by prefixing the vectors with the collection id and the name of the vector index path (we call this the Flat index). This can be done easily in the user request path as part of the ”Document Analysis” during ingestion on the primary replica, and then replicated to all secondaries. A range scan in the Bw-Tree for the appropriate prefix would need many fewer random accesses for smaller vectors. This is still not a great option since vectors tend to be large and storing them twice limits scale up.
[…]
However, an exhaustive scan in quantized space does not scale to larger replicas. Cosmos DB can fit over 10 million 768-dimensional floating point vectors in one replica, and we need an index that can answer queries by accessing fewer nodes in Bw-Tree. We use the $\rm DiskANN$ graph-structured index for this. With the $\rm DiskANN$ vector indexing policy in a collection, two additional indexing terms in the Bw-Tree are created: quantized vectors (akin to Q-Flat) and graph adjacency list terms which represent out-neighbors of each vector in the index (see Fig. 2). In the rest of this section, we describe how we layer these index terms in Cosmos DB, and how $\rm DiskANN$ manipulates the terms to update and query the index.
@@ 3.1. Re-designing DiskANN for databases
$\rm DiskANN$ was previously written as a library that manages its own buffers and index layout, thus limiting its use inside a database. We rewrote it in Rust using the following principles so that it can be used with many systems including databases and key-value stores.
@@
The index layout is not controlled or even visible to the algorithms. The core of the library consists of methods that update and query the index by reading and updating index terms – quantized vectors, full precision vectors and neighbor list for vertices – via implementations of Provider traits. The NeighborProvider trait, for example, defines the way the index would retrieve, append and overwrite the out-neighbors of a vector. These traits are implemented by the database which specifies the layout and encoding/decoding for each kind of indexing terms. The database also owns the persistence and recovery of these terms.
@@
The term that the index needs may be immediately available in a memory buffer or may need to be retrieved from a slower storage device. Therefore, the Provider traits allow for get methods to return either the actual data or a future that eventually returns the data so that the calling thread can be scheduled with other work meanwhile. This is encoded via the MaybeDone enumerated type in the snippet below:
@@
Consequently, all update and query methods in the library are also asynchronous and need a runtime to manage threads and drive the futures to completion. We use the Tokio runtime (tok, 2016).
@@
Our rewrite achieved these goals without compromising on performance compared to monolithic “in-memory” libraries such as (et al., 2019; Simhadri et al., 2023). We can implement the Provider traits using a type backed by memory buffers for maximum performance – in fact, the new library is at least as fast as the previous monolithic $\rm DiskANN$ library for all use cases it supported.